OpenStudy (abbycross167):

Can someone please help me answer this math question? Please? I'll fan and give a medal!! I'm working on polynomial division Question - x^2(5x+6) – 3x(5x+6) ------------------- 5x+6

OpenStudy (abbycross167):

@freckles do you by chance know how to do this kind of math?

OpenStudy (ebayminer126):

hmm

OpenStudy (freckles):

both terms in the numerator have a common factor of (5x+6)

OpenStudy (freckles):

the denominator also shares that factor

OpenStudy (freckles):

divide top and bottom by (5x+6)

OpenStudy (abbycross167):

so x^2(5x+6) --------- 5x+6 = x^2?

OpenStudy (freckles):

\[\frac{x^2(5x+6)-3x(5x+6)}{5x+6} \\ =\frac{x^2(5x+6)}{5x+6}-\frac{3x(5x+6)}{5x+6}\] yes the first term simplifies to x^2

OpenStudy (freckles):

what about the second term

OpenStudy (abbycross167):

And 3x(5x+6) ------- 5x+6 = 3x?

OpenStudy (freckles):

and you make sure you put a minus between those terms since there is a minus in between them in the numerator

OpenStudy (freckles):

\[\frac{x^2(5x+6)-3x(5x+6)}{5x+6} \\ =\frac{x^2(5x+6)}{5x+6}-\frac{3x(5x+6)}{5x+6} \\ =x^2-3x\]

OpenStudy (abbycross167):

ok awesome! that's what I had written as the steps! thank you so much!! I have one more problem to do, would you mind checking my steps to that problem?